# Exploring Persona Sentiment Sensitivity in Personalized Dialogue Generation

### v1

@@ Abstract
Personalized dialogue systems have advanced considerably with the integration of user-specific personas into large language models (LLMs). However, while LLMs can effectively generate personalized responses, the influence of persona sentiment on dialogue quality remains underexplored. In this work, we conduct a large-scale analysis of dialogues generated using a range of polarized user profiles. Our experiments reveal that dialogues involving negatively polarized users tend to overemphasize persona attributes, leading to increased entailment and contradiction instances and lower overall coherence. In contrast, positively polarized profiles yield dialogues that selectively incorporate persona information, resulting in smoother and more coherent interactions. Furthermore, we find that personas with weak or neutral sentiment generally produce lower-quality dialogues. Motivated by these findings, we propose a dialogue generation approach that explicitly accounts for persona polarity by combining a turn-based generation strategy with a profile ordering mechanism. Our study provides new insights into the sensitivity of LLMs to persona sentiment and offers guidance for developing more robust and nuanced personalized dialogue systems.
@@ 1 Introduction
Personalized dialogue systems have evolved through advances in dataset construction Li et al. (2016); Zhang (2018), training methodologies Kingma (2013); Bowman et al. (2015), and studies on integrating user-specific preferences to enhance conversational abilities Zhang (2019); Roller (2020). With the advent of Large Language Models (LLMs), embedding personas via system prompts Yang et al. (2023); Wang et al. (2023b) has become increasingly effective. This progress has enabled applications ranging from personalized chatbots Shuster et al. (2022); Lee et al. (2023) to enhanced performance in downstream tasks through inter-LLM conversations Park et al. (2023) and synthetic dataset generation Jandaghi et al. (2023). These developments underscore the growing importance of personalization in dialogue systems.
@@
However, despite these advances, recent studies have shown that LLMs are highly sensitive to contextual sentiment polarity Liu et al. (2024); Wu et al. (2024b). In contrast, the impact of assigned persona sentiment on dialogue quality remains less explored. For example, Figure 1 illustrates dialogues generated by users with distinctly positive and negative personas. In dialogues driven by positive personas (shown in blue), the integration of persona attributes occurs in a manner that supports coherence and aligns naturally with the flow of conversation. Conversely, dialogues featuring negative personas (shown in red) often display contradictions and inconsistencies, as the persona attributes may be overemphasized or misaligned with the dialogue context. These findings highlight the necessity of systematically examining how variations in persona polarity influence LLM behavior.
In this paper, we construct and analyze personalized dialogues generated across a spectrum of polarized profiles. (§3) Our extensive analysis of personalized dialogues generated across a diverse spectrum of polarized profiles reveals several key insights: 1) Dialogues between negatively polarized users tend to overemphasize persona attributes, which leads to increased contradictions and reduced overall coherence. 2) Dialogues between positively polarized users selectively incorporate persona information, resulting in interactions that are both coherent and contextually appropriate. 3) Profiles with weak or neutral sentiment generally yield lower-quality dialogues, as evidenced by significant differences in consistency and coherence metrics.
These insights underscore the need for methods that explicitly account for persona polarity in the design of personalized dialogue systems. To address this challenge, we propose a dialogue generation approach that combines a turn-based strategy with a profile ordering mechanism. By positioning personas containing neutral, or weak sentiment earlier and positive personas later in the dialogue, this approach aims to achieve more consistent and coherent interactions (§4). Our findings shed light on LLM sensitivity to persona sentiment and offer valuable guidance for developing more nuanced and effective personalized dialogue systems.
@@ 2 Study Design and Experimental Setup
This section explains the overall basic setup for our study. We first introduce the dataset and describe how we construct polarized user profiles. Next, we present our research questions, detail the dialogue generation strategies, and describe the evaluation metrics used to assess dialogue quality. We use four RTX 4090 GPUs to generate and evaluate all the dialogues.
@@ 2.2 Polarized User Profile Construction
Each persona in the ConvAI2 training set spans various content domains and carries a distinct sentiment polarity. We use the distilbert-base-uncased-finetuned-sst-2-english model HF Canonical Model Maintainers (2022) to classify the sentiment of each sentence, retaining only those with a confidence score above 0.99 as polarized personas. Consequently, as shown in Figure 2, 2,691 are classified as positive, 1,006 as negative, and 2,429 as neutral (weak), indicating that polarized personas constitute the majority (around 60% of all personas). Notably, approximately 17% of the dataset’s personas are negatively polarized—a higher proportion than one might intuitively expect. These figures suggest that polarized personas are not merely reflect a contrived scenario fabricated for the experiment. For detailed statistics, see Appendix D.
@@
To create a user profile, we combine $K$ polarized personas, where $K$ is the number of personas we include in each profile. To avoid internal contradictions, we follow the procedure in Jandaghi et al. (2023): starting with an initial persona, we iteratively add additional personas only if the nli-deberta-v3-large model He et al. (2021) detects no conflicts with the existing set. This approach yields three profile types:
@@ 2.4 Dialogue Generation Strategy
Traditional personalized dialogue systems use a turn-based approach to generate responses sequentially based on a given persona and dialogue context Zhou et al. (2023); Tseng et al. (2024). However, recent advances in LLMs have enabled a dual-persona joint generation method that encodes both personas simultaneously to generate the entire dialogue in one pass Liu et al. (2020); Xu et al. (2022); Jandaghi et al. (2023). For resource efficiency, we primarily use dual-persona joint generation for large-scale dialogue production and analysis, while also experimenting with turn-based generation to identify optimal strategies (§4. Both methods use greedy decoding (zero-temperature), with a maximum length of 4096 tokens for dual-persona joint generation and 128 tokens for turn-based generation. Prompt templates for both strategies are provided in Appendix C. After generation, we filter out outliers, such as refusal-to-answer responses or repeated utterances inside a dialogue.
@@ 2.5 Evaluation Metrics
We adopt the C score Madotto et al. (2019), a common metric in personalized dialogue systems that calculates entailment scores for user persona-utterance pairs using an NLI model. Detailed adaptations of the C score for our task are provided in Appendix B.1. Since summing raw scores may lead to Simpson’s Paradox Simpson (1951) due to varying entailment and contradiction counts, we introduce the Contradiction Ratio (Contd.), which measures the proportion of contradictions among all cases. Additionally, we propose a novel metric called the perplexity gap (P gap). Using GPT2-large Radford et al. (2019), we compute the difference between the perplexity of a dialogue $D$ and the perplexity of $D$ conditioned on two user profiles $U_{1}$ and $U_{2}$, quantifying how much the user profiles influence dialogue generation.
@@
We assess dialogue coherence using the perplexity (Perp.) of the dialogue Song et al. (2021); Zhou et al. (2023), estimated with GPT2-large Radford et al. (2019). Additionally, we employ two automated coherence evaluation frameworks: Q-DCE Ye et al. (2021), which uses a BERT-based encoder to predict a coherence score, and PairEval Park et al. (2024), which utilizes a fine-tuned LLaMA-2 model to compare generated dialogues with reference responses. Implementation details for these metrics are provided in Appendix B.2.
@@
Recent studies have shown that LLM-based evaluations can serve as a cost-effective alternative to human annotation Chiang and Lee (2023). Inspired by Chain-of-Thought prompting for summarization Liu et al. (2023), we adapt a similar approach with GPT-4o Hurst et al. (2024) to evaluate dialogue quality, following  Jandaghi et al. (2023). Our prompt format, along with explicit guidelines for each evaluation dimension, allows the model to provide consistent judgments. Full prompt details are available in Appendix C.
Method Pairing Consistency Coherence
@@ 3.1 Does Dialogue Quality Diverge According to Polarized User-Pairing?
Inspired by psychological research indicating that the quality of conversations varies with participants’ traits Hassan et al. (2019), this section explores how polarized profiles and pairing types affect dialogue outcomes under various configurations. We conduct a large-scale, detailed study employing four backbone LLMs and eight metrics to verify result robustness across models and evaluation measures.
@@ 3.1.1 Experiment Setup
For each pairing configuration, we feed all user pairs into four modern LLMs—LLaMa-3.1-8B-Instruct Dubey et al. (2024), Qwen-2.5-7B-Instruct Yang et al. (2024), Ministral-8B-Instruct Jiang et al. (2024), and Gemma-2-9B-Instruct Team et al. (2024)—which collectively generate 60K dialogues and filter out outliers, as mentioned in §2.4. We ensure that every model processes the same user pairs in the identical order. Dialogues where the model refuses to respond or produces repetitive utterances are discarded, resulting in approximately 58K valid dialogues. We provide a detailed distribution of dialogues by pairing scenario and model in Appendix D. Finally, we apply four consistency metrics and four coherence metrics to all dialogues, aggregating the results by pairing type and LLM.
@@ 3.1.2 Results & Analysis
Positive-Polarized Configurations Yield Significantly Better Dialogues
@@ 3.2.1 Experimental Setup
We aim to determine whether dialogues between users with strongly polarized personas (e.g., clearly positive or negative) differ from those with weak or neutral emotional content. We define each persona’s polarity level based on the confidence score produced by the classification model used in §2.2: values closer to 0 indicate a highly negative polarity, those closer to 1 indicate a highly positive polarity, and scores near 0.5 represent a neutral persona. Because half of all personas lie at sentimental extremes, uniformly dividing confidence scores would yield imbalanced distributions across levels. Instead, we partition the score range into nine intervals that ensure a more balanced representation of persona instances. Further details on the partition criteria and the distribution across levels are provided in Appendix D.
@@ 4 How to Make LLMs Robust to Polarity?
Table 2: Dialogue consistency and coherence across generation strategies using LLaMa-3.2-3B and Qwen-2.5-7B with the original pairing. Abbreviations: asc. (score ascending), dsc. (score descending), c-asc. (center-out score ascending). Best results are in bold.
4.1 Adjusting Generation Strategy
Additionally, generating utterances on a turn-by-turn basis sometimes results in excessively long responses that cluttered the overall dialogue. To address this, we add an instruction in the prompt to produce short and concise utterances. The detailed prompt is provided in Appendix C.
4.2 User Profile Ordering
[…]
We apply this ordering only to the turn-based agent conversation method, as the joint generation method combines both profiles into a single prompt.
4.3 Results & Analysis
@@
For the LLaMa-3.2-3B model, the turn-based approach significantly enhances dialogue consistency compared to dual-persona joint generation. This improvement stems from separating users’ polarized personas into distinct inputs, thereby reducing internal complexity.
In contrast, the Qwen-2.5-3B model demonstrates lower G-eval scores primarily because it blends the personas of both users into a single dialogue. When User 1 initiates the conversation, its persona tends to dominate, leading User 2 to adopt a similar framing. For example, if User 1’s persona is "I enjoy sewing", the dialogue might begin with, "I heard you’re good at sewing", prompting User 2 to continue discussing sewing. As a result, User 2’s responses inadvertently mix in User 1’s persona, while User 1 seldom reaffirms the persona initially designated for User 2. This behavior ultimately lowers the overall consistency score.
Nonetheless, contradictions between personas and utterances actually decrease, boosting contradiction metrics (C score and Contd.). Also, we can resolve this issue by prompt engineering, which can be further explored.

### v2

@@ Abstract
Personalized dialogue systems have advanced considerably with the integration of user-specific personas into large language models (LLMs). However, while LLMs can effectively generate personalized responses, the influence of persona sentiment on dialogue quality remains underexplored. In this work, we conduct a large-scale analysis of dialogues generated using a range of polarized user profiles. Our experiments reveal that dialogues involving negatively polarized users tend to overemphasize persona attributes. In contrast, positively polarized profiles yield dialogues that selectively incorporate persona information, resulting in smoother interactions. Furthermore, we find that personas with weak or neutral sentiment generally produce lower-quality dialogues. Motivated by these findings, we propose a dialogue generation approach that explicitly accounts for persona polarity by combining a turn-based generation strategy with a profile ordering mechanism and sentiment-aware prompting. Our study provides new insights into the sensitivity of LLMs to persona sentiment and offers guidance for developing more robust and nuanced personalized dialogue systems.***Our implementation is publicly available at https://github.com/imsongpasimin/PesonaSensitivity
@@ 1 Introduction
Personalized dialogue systems have evolved through advances in dataset construction Li et al. (2016); Zhang (2018), training methodologies Kingma (2013); Bowman et al. (2015), and studies on integrating user-specific preferences to enhance conversational abilities Zhang (2019); Roller (2020). With the advent of Large Language Models (LLMs), embedding personas via system prompts Yang et al. (2023); Wang et al. (2023b) has become effective. This progress has enabled applications ranging from personalized chatbots Shuster et al. (2022); Lee et al. (2023) to enhanced performance in downstream tasks through inter-LLM conversations Park et al. (2023) and synthetic dataset generation Jandaghi et al. (2023). These developments underscore the growing importance of personalization in dialogue systems.
@@
However, despite these advances, recent studies have shown that LLMs are highly sensitive to contextual sentiment polarity Liu et al. (2024); Wu et al. (2024b). In contrast, the impact of assigned persona sentiment on dialogue quality remains less explored. For example, Figure 1 illustrates dialogues generated by users with distinctly positive and negative personas. In dialogues driven by positive personas, the integration of persona attributes occurs in a manner that supports coherence and aligns naturally with the flow of conversation. Conversely, dialogues featuring negative personas often display contradictions and inconsistencies, as the persona attributes may be overemphasized or misaligned with the dialogue context. These findings highlight the necessity of systematically examining how variations in persona polarity influence LLM behavior.
In this paper, we construct and analyze personalized dialogues generated across a spectrum of polarized profiles. (§3) Our extensive analysis of personalized dialogues generated across a diverse spectrum of polarized profiles reveals several key insights: 1) Dialogues between negatively polarized users tend to overemphasize persona attributes, which leads to increased contradictions and reduced overall coherence. 2) Dialogues between positively polarized users selectively incorporate persona information, resulting in interactions that are both coherent and contextually appropriate. 3) Profiles with weak or neutral sentiment generally yield lower-quality dialogues, as evidenced by significant differences in various metrics.
These insights underscore the need for methods that explicitly account for persona polarity in the design of personalized dialogue systems. To address this challenge, we propose a dialogue generation framework that combines a turn-based strategy, a profile ordering mechanism, and a sentiment-aware prompting. Specifically, personas exhibiting neutral or weak sentiment are placed earlier in the conversation, whereas more positive personas are positioned later. A concise prompt further instructs the model to attend carefully to negative or neutral sentiments when generating responses. These approaches promote more consistent and coherent interactions (§4). Our findings shed light on LLM sensitivity to persona sentiment and offer valuable guidance for developing more nuanced and effective personalized dialogue systems.
@@ 2 Study Design and Experimental Setup
This section explains the overall basic setup for our study. We first introduce the dataset and describe how we construct polarized user profiles. Next, we present our research questions, detail the dialogue generation strategies, and describe the metrics used to assess the dialogues. We use four RTX 4090 GPUs to generate and evaluate all the dialogues.
@@ 2.2 Polarized User Profile Construction
Each persona in the ConvAI2 spans various content domains and carries a distinct sentiment polarity. We use the distilbert-base-uncased-finetuned-sst-2-english model HF Canonical Model Maintainers (2022) to classify the sentiment of each sentence, retaining only those with a confidence score above 0.99 as polarized personas. Consequently, as shown in Figure 2, 2,691 are classified as positive, 1,006 as negative, and 2,429 as neutral (weak), indicating that polarized personas constitute the majority (around 60% of all personas). Notably, approximately 17% of the dataset’s personas are negatively polarized—a higher proportion than one might intuitively expect. These figures suggest that polarized personas are not merely reflect a contrived scenario fabricated for the experiment. For detailed statistics, see Appendix F.
@@
To create a user profile, we combine $K$ polarized personas, where $K$ is the number of personas we include in each profile. To avoid internal contradictions, we follow the procedure in Jandaghi et al. (2023): starting with an initial persona, we iteratively add additional personas only if the nli-deberta-v3-large He et al. (2021) detects no conflicts with the existing set. This approach yields three profile types:
@@ 2.4 Dialogue Generation Strategy
Traditional personalized dialogue systems use a turn-based approach to generate responses sequentially based on a given persona and dialogue context Zhou et al. (2023); Tseng et al. (2024). However, recent advances in LLMs have enabled a dual-persona joint generation method that encodes both personas simultaneously to generate the entire dialogue in one pass Liu et al. (2020); Xu et al. (2022); Jandaghi et al. (2023). For resource efficiency, we primarily use dual-persona joint generation for large-scale dialogue production and analysis, while also experimenting with turn-based generation to identify optimal strategies (§4. Both methods use greedy decoding (zero-temperature), with a maximum length of 4096 tokens for dual-persona joint generation and 128 tokens for turn-based generation. Prompt templates for both strategies are provided in Appendix E. After generation, we filter out outliers, such as refusal-to-answer responses or repeated utterances inside a dialogue.
@@ 2.5 Evaluation Metrics
We adopt the C score Madotto et al. (2019), a common metric in personalized dialogue systems that calculates entailment scores for user persona-utterance pairs using an NLI model. Detailed adaptations of the C score for our task are provided in Appendix D.1. Since summing raw scores may lead to Simpson’s Paradox Simpson (1951) due to varying entailment and contradiction counts, we introduce the Contradiction Ratio (Contd.), which measures the proportion of contradictions among all cases. Additionally, we propose a novel metric called the perplexity gap (P gap). Using GPT2-large Radford et al. (2019), we compute the difference between the perplexity of a dialogue $D$ and the perplexity of $D$ conditioned on two user profiles $U_{1}$ and $U_{2}$, quantifying how much the user profiles influence dialogue generation.
@@
We assess dialogue coherence using the perplexity (Perp.) of the dialogue Song et al. (2021); Zhou et al. (2023), estimated with GPT2-large Radford et al. (2019). Additionally, we employ two automated coherence evaluation frameworks: Q-DCE Ye et al. (2021), which uses a BERT-based encoder to predict a coherence score, and PairEval Park et al. (2024), which utilizes a fine-tuned LLaMA-2 model to compare generated dialogues with reference responses. Details for these metrics are provided in Appendix D.2.
@@
Recent studies have shown that LLM-based evaluations can serve as an effective alternative to human annotation Chiang and Lee (2023). Inspired by CoT evaluation framework for summarization tasks Liu et al. (2023), we adapt a similar approach with GPT-4o Hurst et al. (2024) to evaluate dialogue quality, following  Jandaghi et al. (2023). Our prompt format, along with explicit guidelines for each evaluation dimension, allows the model to provide consistent judgments. Full prompt details are available in Appendix E.
Human Evaluation
Beyond the automated metrics, we also conduct a human evaluation on a subset of samples. For each configuration under evaluation, we randomly sample 40 persona–dialogue pairs generated by Qwen. Three fluent English speakers then rate these samples using the same criteria as the LLM-based automatic evaluation, but on a condensed 1–3 scale. We set the final score for each configuration to the average of the three evaluators’ ratings.
Model Pairing Consistency Coherence
@@ 3.1 Does Dialogue Quality Diverge According to Polarized User-Pairing?
Inspired by psychological research indicating that the quality of conversations varies with participants’ traits Hassan et al. (2019), this section explores how polarized profiles and pairing types affect dialogue outcomes under various configurations. We conduct a large-scale, detailed study employing four backbone LLMs and eight metrics to verify the robustness of the results.
@@ 3.1.1 Experiment Setup
For each pairing configuration, we feed all user pairs into four modern LLMs—LLaMa-3.1-8B-Instruct Dubey et al. (2024), Qwen-2.5-7B-Instruct Yang et al. (2024), Ministral-8B-Instruct Jiang et al. (2024), and Gemma-2-9B-Instruct Team et al. (2024)—which collectively generate 60K dialogues and filter out outliers, as mentioned in §2.4. We ensure that every model processes the same user pairs in the identical order. Dialogues where the model refuses to respond or produces repetitive utterances are discarded, resulting in approximately 58K valid dialogues. We provide a detailed distribution of dialogues by pairing scenario and model in Appendix F. Finally, we apply four consistency metrics and four coherence metrics to all dialogues, aggregating the results by pairing type and LLM. For utilizing GPT-4o as backbone LLM, we report the results in Appendix C.
@@ 3.1.2 Results & Analysis
Positive Configurations Yield Better Dialogues
@@
Human Evaluation
We verify whether the trends identified by automatic metrics persist under human judgment by evaluating dialogues generated with the Original pairing, Negative pairing, and Positive pairing configurations.
Pairing Consistency Coherence
Original 2.36 2.01
Negative 2.40 2.12
Positive 2.51 2.30
Table 2: Human evaluation results on consistency and coherence among the sentiment of user profiles.
As shown in Table 2, human raters corroborate the automatic findings: dialogues generated from polarized users—especially Positive pairings—exhibit the highest quality. Notably, the decline in scores for the Original pairing is far more pronounced in the human evaluation than in the automatic assessment, further confirming the validity of our observations.
@@ 3.2.1 Experimental Setup
We aim to determine whether dialogues between users with strongly polarized personas (e.g., clearly positive or negative) differ from those with weak or neutral emotional content. We define each persona’s polarity level based on the confidence score produced by the classification model used in §2.2: values closer to 0 indicate a highly negative polarity, those closer to 1 indicate a highly positive polarity, and scores near 0.5 represent a neutral persona. Because half of all personas lie at sentimental extremes, uniformly dividing confidence scores would yield imbalanced distributions across levels. Instead, we partition the score range into nine intervals that ensure a more balanced representation of persona instances. Further details on the partition criteria and the distribution across levels are provided in Appendix F.
@@ 3.2.2 Results & Analysis
We further analyze the underlying causes of the performance gap discussed in this section in Appendix A.
@@ 4 How to Make LLMs Robust to Polarity?
4.1 Adjusting Generation Strategy
Additionally, generating utterances on a turn-by-turn basis sometimes results in excessively long responses that cluttered the overall dialogue. To address this, we add an instruction in the prompt to produce short and concise utterances. The detailed prompt is provided in Appendix E.
4.2 User Profile Ordering & Sentiment-aware Prompting
[…]
We apply this ordering only to the turn-based approach, as the joint generation method combines both profiles into a single prompt.
Plus, we concatenate a lightweight prompt template: "Please ensure that each user’s persona, especially negative or neutral personas, is well integrated into the dialogue and that the overall dialogue remains coherent". It selectively conditions the model on the sentiment of personas present in a user profile at the end of the generation prompt. Because large language models already handle positive personas reliably, the template explicitly instructs the model to be aware when responding to negative or neutral personas.
4.3 Results & Analysis
@@
+ sap. 0.777 8.27 4.58 2.61 4.63
+ c-sap. 0.774 7.49 4.59 2.69 4.77
Table 3: Dialogue consistency and coherence across generation strategies using LLaMa-3.2-3B and Qwen-2.5-7B with the original pairing. Abbreviations: asc. (score ascending), dsc. (score descending), c-asc. (center-out score ascending), sap. (sentiment-aware prompting), c-sap. (center-out score ascending + sentiment-aware prompting). Best results are in bold.
@@
For the LLaMa-3.2-3B model, the turn-based approach significantly enhances dialogue consistency compared to dual-persona joint generation. This improvement stems from separating users’ polarized personas into distinct inputs, thereby reducing internal complexity. In contrast, the Qwen-2.5-3B model demonstrates lower G-eval scores primarily because it blends the personas of both users into a single dialogue. When User 1 initiates the conversation, its persona tends to dominate, leading User 2 to adopt a similar framing. For example, if User 1’s persona is "I enjoy sewing", the dialogue might begin with, "I heard you’re good at sewing", prompting User 2 to continue discussing sewing. As a result, User 2’s responses inadvertently mix in User 1’s persona, while User 1 seldom reaffirms the persona initially designated for User 2. This behavior ultimately lowers the overall consistency score. Nonetheless, contradictions between personas and utterances actually decrease, boosting contradiction metrics (C score and Contd.). Also, we can resolve this issue by prompt engineering, which can be further explored.
@@
Synergy between Profile Ordering and Sentiment-Aware Prompting
Furthermore, as shown in Table 3, adding the sentiment-aware prompt alone produces some improvements for both models. More importantly, jointly applying the center-out score-ascending strategy and sentiment-aware prompting yields a strong synergy that markedly elevates dialogue quality. In particular, the coherence metric increases by around 0.1 points for both models, demonstrating substantial gains in conversational coherence. We attribute this improvement to positioning the personas that require heightened attention earlier in the prompt, thereby capitalizing on the uni-directional decoding of large language models in concert with the sentiment-aware instruction.
Human Evaluation
We further manually compare turn-based generation with and without our profile ordering and sentiment-aware prompting techniques.
Strategy Consistency Coherence
Turn-based 1.80 2.40
+ c-sap. 2.27 (+0.47) 2.43 (+0.03)
Table 4: Human evaluation results for consistency and coherence of dialogues generated by the turn-based approach, with and without profile ordering and sentiment-aware prompting.
As shown in Table 4, under the original pairing condition, applying our profile ordering and sentiment-aware prompting techniques to the turn-based generation improves consistency by 0.47 points and coherence by 0.03 points compared with the baseline. These observations further reinforce the robustness and efficacy of our proposed approach.